OpenStudy (anonymous):

Find the first partial derivatives of f(x,y) = [(2x − 2y)/(2x +2y)] at the point (x,y) = (3, 1). [(∂f)/(∂x)](3, 1) = ??? [(∂f)/(∂y)](3, 1) = ???

OpenStudy (turingtest):

w/respect to y: [-2(2x+2y)-2(2x-2y)]/(2x+2y)^2 =-4x/(2x+2y)^2=-x/(x+y)^2 evaluated at (3,1) -3/16

OpenStudy (turingtest):

partial w/respect to x: [2(2x+2y)-2(2x-2y )]/(2x+2y)^2 =4y/(2x+2y)^2=y/(x+y)^2 evaluated at (3,1)=1/16

OpenStudy (turingtest):

sorry i forgot to square the bottom the first time. my bad...